OpenStudy (anonymous):

While building pyramids, ancient Egyptians used levers to help lift stones weighing several tons. The amount of work, W, can be calculated with the equation W = Fd, where F is the force to the stones and d is the distance the stones were moved. Which equation can be used to solve for distance?

OpenStudy (anonymous):

solve for the letter d :) *divide the f over to the other side to isolate d \[{W\over f }= d\]or\[d={W\over f }\]

OpenStudy (anonymous):

i hope this helps! @krazyman :)

OpenStudy (anonymous):

so d divided by f

OpenStudy (anonymous):

no its like this \[{df\over f}={W\over f}\]

OpenStudy (anonymous):

so the f cancels out and leaves the d :)

OpenStudy (anonymous):

do you get it? :)

OpenStudy (anonymous):

a d= W F b d = W – F c d = F – W d d = W ÷ F the answers which one

OpenStudy (anonymous):

\[d={W\over f}\] :)

OpenStudy (anonymous):

is that the first second third or fourth

OpenStudy (anonymous):

\[{W \over f}=W \div f\]

OpenStudy (anonymous):

the answer is D...